OpenStudy (anonymous):

A square has sides x cm long. A kite has one diagonal x cm long. Find the length of the second diagonal for the kite that makes its area the same as the area of the square.

OpenStudy (raden):

1/2 * x * d = x^2 solve for d

OpenStudy (anonymous):

d=1/2x*x^2?

OpenStudy (raden):

it should be d = 2x^2/x = 2x

OpenStudy (anonymous):

is that the answer?

OpenStudy (raden):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

you're welcome